# “When to Hand Off, When to Work Together”: Expanding Human-Agent Co-Creative Collaboration through Concurrent Interaction

### 06 Mar 2026 v2

@@ Abstract.
Human collaborators coordinate dynamically through process visibility and workspace awareness, yet AI agents typically either provide only final outputs or expose read-only execution processes (e.g., planning, reasoning) without interpreting concurrent user actions on shared artifacts. Building on mixed-initiative interaction principles, we explore whether agents can achieve collaborative context awareness—interpreting concurrent user actions on shared artifacts and adapting in real-time. Study 1 (N=10 professional designers) revealed that process visibility enabled reasoning about agent actions but exposed conflicts when agents could not distinguish feedback from independent work. We developed Cleo, which interprets collaborative intent and adapts in real-time. Study 2 (N=10, two-day with stimulated recall interviews) analyzed 214 turns, identifying five action patterns, six triggers, and four enabling factors explaining when designers choose delegation (70.1%), direction (28.5%), or concurrent work (31.8%). We present a decision model with six interaction loops, design implications, and an annotated dataset.
Human-Agent Collaboration, Interaction Pattern Analysis, Concurrent Human-Agent Interaction, Generative AI, Co-Creative Design Agent
@@ 1. Introduction
In human–human collaborative creative work, as commonly observed in shared tools such as Figma [^1] https://www.figma.com or Google Slides [^2] https://docs.google.com/presentation, visibility into a collaborator’s actions and awareness of who is working on what are fundamental for effective coordination (Dourish and Bellotti, 1992). Collaborators observe each other’s ongoing work, anticipate next steps, intervene at opportune moments, and fluidly divide tasks based on emerging priorities and expertise (Tang, 1991). This dynamic coordination relies on process transparency—the ability to see how work unfolds, not just what is produced—and mutual awareness of presence and focus within the shared workspace (Gutwin and Greenberg, 2002; Gutwin et al., 1996). These elements enable richer collaboration patterns beyond sequential turn-taking, including concurrent work, interleaved workflows, and real-time adaptation to each other’s contributions (Tang, 1991).
AI agents increasingly perform tasks step-by-step using human-interpretable operations (e.g., checking files, clicking buttons), making their processes observable across domains like programming (Sapkota et al., 2025), web automation (Jiang and Xia, 2025), and research (Gottweis et al., 2025). In creative domains, however, agents typically provide only final outputs rather than exposing execution processes (Make, ; Stitch, ; Lovable, ), limiting them to assistant roles for specific subtasks (You et al., 2025; Zhou et al., 2024). Yet process observability alone is insufficient. The conceptual challenge of flexibly shifting between delegation and direct control based on context has deep roots in mixed-initiative interaction (Horvitz, 1999) and adjustable autonomy in robotics (Pellegrinelli et al., 2016). What remains unexplored is whether agents in creative contexts can go beyond read-only process transparency to achieve collaborative context awareness—interpreting concurrent user direct manipulation on shared artifacts and adapting execution in real-time, enabling the fluid, interleaved collaboration patterns seen in human–human creative work.
We investigate this question through two design probes in Figma. Our first probe established process transparency through voice interaction, visual tool operations, and workspace awareness (agent cursor presence). Study 1 (N=10 professional designers) revealed that while step-by-step execution visibility enabled anticipation and timely intervention, designers expected mid-execution concurrent collaboration—wanting to interleave their work with the agent’s ongoing execution. The agent, however, could not distinguish user feedback from independent work, creating conflicts that revealed a critical gap: the absence of collaborative context awareness.
@@
This finding motivated Cleo (Collaborative Linked Executive Operator), which tracks concurrent user actions, interprets collaborative intent (feedback vs. independent work), and adapts in real-time. Study 2 (N=10, two-day exploratory study with stimulated recall interviews (Dempsey, 2010)) analyzed 214 turn-level interactions to understand when and why designers choose different collaboration modes. Through qualitative coding, we identified five action patterns (hands-off, observational, concurrent, directive, terminating in Figure 7), six triggers prompting intervention (e.g., idea spark, task mismatch in Figure 8), and four enabling factors determining choices (e.g., mental model, task priority in Figure 9). Concurrent work appeared in 31.8% of turns, full delegation in 70.1%, and directive intervention in 28.5%, with dynamic shifts based on evolving priorities.
We synthesized these findings into a decision model (Figure 10) that characterizes how designers dynamically navigate among five actions: delegation, observation, direction, concurrent work, and termination. The model identifies six interaction loops driven by triggers, enabling factors, and evolving priorities. Drawing from the dynamics revealed by these loops, we derive design implications for adaptive collaborative agents that support seamless mode transitions and context-aware adaptation. We also release our dataset of 214 interaction turns with logged actions, triggers, and designer rationales to support future research in human-agent co-creative collaboration.
Our contributions include:
@@
  • A characterization of concurrent human-agent interaction with five user action patterns enabling fluid collaboration during agent execution.
@@
  • Design implications for adaptive collaborative agents supporting seamless mode transitions and context-aware adaptation.
  • A dataset of 214 human-agent co-creative interaction turns with logged actions.
@@ 2.1. Dynamics of Flexible Collaboration: From Turn-taking to Synchrony
Human collaboration is a dynamic spectrum shifting between different interaction modes (Tang, 1991). While turn-taking facilitates coordination in complex tasks by minimizing cognitive load and conflicts through sequential actions (Kirschner et al., 2009; Sacks et al., 1974), creative domains often demand synchronous collaboration (Nam and Sakong, 2009). This is because creative work is rarely a linear process of executing predefined steps; rather, it involves the co-evolution of problem and solution, where design ideas are refined through immediate visual feedback and spontaneous mutual inspiration (Maher and Poon, 1996; Ishii et al., 1993). In such contexts, waiting for a turn can break the creative flow and delay the rapid synthesis of ideas. Therefore, partners often bump in for immediate adjustments or perform simultaneous edits to interleave workflows, ensuring that the evolving design remains a shared mental model (Dourish and Bellotti, 1992; Tang, 1991).
@@
Modern tools like Figma and Google Slides support this fluidity by fostering workspace awareness—the real-time understanding of a collaborator’s presence and intentions. This is enabled through visual cues like telepointers (moving cursors) and object highlighting, allowing users to anticipate moves and intervene without explicit communication (Gutwin and Greenberg, 2002; Gutwin et al., 1996). However, most co-creative AI agents remain confined to a rigid wait-your-turn paradigm. Unlike the parallel dynamics mastered in human-human work, human-AI interaction is still largely limited to a prompt-and-output cycle, lacking the mutual awareness essential for true creative partnership.
$2.$ Paradigm Shifts in Human-Agent Collaboration
@@
AI is evolving from a command-based tool into a proactive “co-worker” that actively participates within shared workspaces (Horvitz, 1999). This shift has been driven by increased AI capabilities and frameworks that enable models to leverage diverse tools (e.g., code, web search, etc.) (Schick et al., 2023) and interleave reasoning with actions across extended time horizons (Yao et al., 2023). These capabilities have enabled diverse agents for programming (by Anthropic, ; Jimenez et al., 2023; Sapkota et al., 2025), web automation (AI, ; Deng et al., 2023; Jiang and Xia, 2025), visual design (Make, ), research (Gottweis et al., 2025), data analysis (Hong et al., 2025) etc. As these agents expose their reasoning and intermediate actions within their generation, this naturally affords greater interpretability by allowing users to monitor the agent behavior in real time (Amershi et al., 2019; Prasongpongchai et al., 2025).
@@
However, despite this increased visibility, most collaboration between human users and AI agents remains rooted in turn-taking. Even when users can interpret what actions the agent is taking on which artifacts, they can rarely act simultaneously on those artifacts. This limits users’ ability to correct agents mid-execution (Amershi et al., 2019) to prevent misalignments as the task progresses. In this work, we explore the potential of concurrent human-agent collaboration not as a replacement for turn-taking, but as an additional mode for collaboration—enabling agents to move beyond delegated executors to responsive partners capable of accommodating spontaneous interventions within a shared workspace.
@@ 2.3. Design Agents: From Workflow Automation to Process-Visible Collaboration
Existing agents for supporting the design process can be categorized into three distinct approaches based on how they structure their tasks and interact with users. The first category is subtask design agents, which support specific stages within an existing design workflow by automating specific and tedious subtasks, such as adding the details (Lawton et al., 2023; Karimi et al., 2020; Zhou et al., 2024) or generating diverse alternatives (Make, ). In this model, the overall workflow remains human-driven, while the agent functioning as an assistive role. The second category is process-structured agents, which operate within a predefined step-by-step design activities through node-based workflow (Hao et al., 2025; Wang et al., 2025; Ding et al., 2023; You et al., 2025). These systems expose intermediate process steps within the whole workflow and support guided co-creation between agent and human, but the collaboration is constrained by predefined logical nodes. The third type is the output-centric agent, which exposes an internal reasoning process for generating a final output to users in text, such as their execution planning (Stitch, ; Lovart, ; Khurana et al., 2025).
Despite these differences in task structuring, these agents share a common interaction characteristic that the agent’s execution process is hidden from user’s workspace. Users interact with agents only through requests and final outputs rather than through ongoing execution steps. As a result, users are required to wait for final results without seeing the work in progress. While agents in other domains (e.g., programming) increasingly expose execution steps through direct tool operations, design tools have yet to explore how “live” tool-use—such as visible cursor movements and object manipulation—could reshape human-agent collaboration. This suggests an open question of whether process visibility could enable more diverse collaboration modes beyond simple turn-taking.
@@ 3. Study 1: Initial Technology Probe Study
To understand how agents’ workflow sharing via human-understandable tool calls affects human-agent collaboration in design, we implemented an initial design probe and conducted an exploratory study with 10 professional designers.
@@ 3.1.1. Three Basic Components for Collaborative Agent Design
We ground the system’s architecture in three fundamental components from human-human collaboration for collaborative agent design: (1) Interaction Modality for peer-like communication, (2) Process Transparency for observable agent work, and (3) Presence Awareness for mutual attention visualization.
Interaction Modality: Voice Interaction with Flexible Object Referencing
The first probe system uses voice-based interaction as the primary channel, enabling natural turn-taking and real-time intention sharing (Hutchins, 1995; Reicherts et al., 2022). Voice creates a sense of horizontal partnership rather than task delegation (Bickmore and Picard, 2005). Furthermore, users can communicate with the system through direct object referencing on the canvas, which enables fluid deictic communication (e.g., “make this larger” while clicking).
Process Transparency: Sharing Implementation Workflow
Unlike general AI design tools that only share final outputs, the system makes its design process visible in real-time through the execution of human design tools (e.g., creating a rectangle to make a button). As the agent generates UI components step by step, users observe elements appearing, positioning, and styling incrementally—ensuring “public availability of actions and artifacts” (Robertson, 2002). This progressive visualization reveals not just what the agent produces, but how it is progressing, establishing common ground around both goals and processes (Cila, 2022; Hill et al., 1992) between the user and the agent.
Presence Awareness: Visualizing Attention in the Shared Workspace
The system displays real-time indicators of agent’s current focus area and working state, following Gutwin and Greenberg’s workspace awareness framework (Gutwin and Greenberg, 2002). The agent character’s location reveals which canvas region the agent is attending to, enabling users to anticipate the next execution area and coordinate accordingly (Dourish and Bellotti, 1992; Fussell et al., 2000). This peripheral awareness creates mutual presence beyond simple loading indicators.
@@ 3.1.2. System Design
Figure 1. Interaction flow of the first probe system. (a) The user clicks the record button to deliver a voice message (b) while selecting canvas elements for context. The transcribed message with selection changes is sent to the agent as input. The agent generates an action plan and iterates through cycles of (1) reason, (2) act, and (3) feedback. (d-f) Each iteration’s progress is reflected immediately on the canvas as tools execute during the act stage, with the agent character floating near modified elements and displaying interim messages.
@@
Interface and Interaction Scenario
The first probe system is implemented as a Figma plugin. Users interact with the system through both the plugin interface and an agent character that appears on the canvas. To provide an input to the agent, users click the record button in the plugin interface to send voice messages (Figure 1A). The system continuously tracks voice input and selection changes on the canvas until the user clicks the stop button. Users can click any element, including reference images or specific frames, to provide accurate context. As illustrated in Figure 1B, when a user says “show color variation of this section” while selecting an element, the transcribed message, along with timestamp and selected element metadata, is sent to the agent as input. Upon receiving user input, the agent responds with a sequence of Figma operations and text messages. The agent completes its response through multiple iterations, executing actions and providing messages incrementally (Figure 1C). For example, rather than displaying style variations all at once, the agent first duplicates the original element (Figure 1D), then applies style changes sequentially (Figure 1E-F), revealing progress step-by-step. At each step, the agent character animates around the modified elements with interim messages, helping users easily track the agent’s actions.
Agent Design and Pipeline
The first probe agent is designed based on the ReAct agent structure (Yao et al., 2023). As illustrated in Figure 1C, it operates through a three-step iteration cycle: (1) reason, (2) act, and (3) feedback. Referring to Figure 2, the cycle begins with the main reasoning stage, which determines whether design actions are required. If the agent deems actions unnecessary, it terminates the cycle and provides a final response through the Message Module. Otherwise, it selects tools from the 38 Figma operations and generates an interim progress message. The selected tools are immediately executed on the canvas alongside the agent character and the interim message.
After each iteration cycle, the summary module generates a text description of what occurred based on the tool execution results. These summaries are retained in context and stored with the chat history. The feedback module then captures the current canvas state and compares it with the previous state before tool execution to assess whether the plan has been fulfilled, and generates a feedback message describing the next steps if needed. If the plan is complete, the agent terminates the loop and proceeds to the message module. Otherwise, it returns to the initial stage of the iteration cycle to begin a new iteration. To prevent infinite loops and ensure usability, we limit the maximum number of iteration cycles to 10.
We utilize Anthropic’s Claude Sonnet 4.5  [^3] https://www.anthropic.com/claude/sonnet as the main reasoning model and Claude Haiku 4.5  [^4] https://www.anthropic.com/claude/haiku for all modules including plan, summary, feedback, and message generation. To enhance reasoning performance, we provide the main model with guidance instructions for Figma, including recommended rules and patterns for UI design. Detailed instructions for each module and the complete set of 38 Figma operation tools can be found in the Appendix. We designed the prompts (e.g., Figma design guideline) by referring Talk To Figma plugin  [^5] https://github.com/grab/cursor-talk-to-figma-mcp.

### 07 Apr 2026 v4

@@ Abstract.
As agents move into shared workspaces and their execution becomes visible, human-agent collaboration faces a fundamental shift from sequential delegation to concurrent co-creation. This raises a new coordination problem: what interaction patterns emerge, and what agent capabilities are required to support them? Study 1 (N=10) revealed that process visibility naturally prompted concurrent intervention, but exposed a critical capability gap: agents lacked the collaborative context awareness needed to distinguish user feedback from independent parallel work. This motivated Cleo, a design probe that embodies this capability, interpreting concurrent user actions as feedback or independent work and adapting execution accordingly. Study 2 (N=10) analyzed 214 turn-level interactions, identifying a taxonomy of five action patterns and ten codes, along with six triggers and four enabling factors explaining when and why users shift between collaboration modes. Concurrent interaction appeared in 31.8% of turns. We present a decision model, design implications, and an annotated dataset, positioning concurrent interaction as what makes delegation work better.
Human-Agent Interaction, Interaction Pattern Analysis, Concurrent Interaction, Collaborative Agent, Creative Workflow
@@ 1. Introduction
AI agents are increasingly capable of executing complex tasks autonomously, such as design (Make, ; Stitch, ; Lovable, ), programming (Sapkota et al., 2025), and web navigation (Jiang and Xia, 2025). As agents move into shared workspaces and their execution becomes visible through emerging tool-calling infrastructures (e.g., Model Context Protocol (Anthropic, 2024)), human-agent collaboration faces a fundamental shift: from sequential delegation, where users specify goals and review outputs, to collaborative co-creation, where users and agents work simultaneously in the same space. This shift raises a new coordination problem: what interaction patterns emerge when humans and agents share a workspace, and what agent capabilities are required to support them?
To explore this, we built a process-transparent agent in Figma that executes design tasks through direct tool operations (e.g., creating frames, applying shadows), helping designers track the agent and coordinate accordingly (Gutwin and Greenberg, 2002; Gutwin et al., 1996; Robertson, 2002). In Study 1 with 10 designers, process visibility revealed how the coordination problem manifests. Designers naturally began intervening mid-execution, interleaving their own work with the agent’s and demonstrating intent through direct manipulation, creating a need for the agent to interpret and respond to these concurrent actions (Figure 1). However, the agent could not interpret these concurrent actions, lacking the collaborative context awareness needed to distinguish feedback to incorporate from independent work to leave alone.
Figure 1. An example of collaborative friction in human-agent collaboration within a shared workspace.
The figure illustrates a conflict scenario in human-agent collaboration on a shared canvas. In the initial state (top), the agent is executing a two-column layout while narrating its progress, and the user concurrently proposes a three-column layout. The circular arrows indicate both parties acting simultaneously on the same workspace. In the resulting state (bottom), the user’s three-column edit is visible on the canvas, but the agent — unaware of the user’s intent — treats it as an error conflicting with its original plan and reverts to the two-column layout. The red spiky border highlights the moment of conflict.
Based on this finding, we developed Cleo (Collaborative Linked Executive Operator), a second design probe that embodies collaborative context awareness by continuously tracking user actions within the workspace, interpreting whether these actions should update the agent’s task planning or be considered as independent work, and adapting the agent’s ongoing execution accordingly (Horvitz, 1999; Pellegrinelli et al., 2016).
@@
With Cleo, we conducted a two-day exploratory study (Study 2, N=10) with stimulated recall interviews (Dempsey, 2010; Calderhead, 1981), analyzing 214 turn-level interactions to understand when and why users choose different collaboration modes. We identified a taxonomy of five action pattern categories and ten codes capturing how users engage with agents during execution (hands-off, observational, concurrent, directive, terminating), six triggers prompting intervention (e.g., idea spark, task mismatch), and four enabling factors shaping those choices (e.g., mental model, task priority). Concurrent work appeared in 31.8% of turns, with dynamic shifts driven by evolving priorities. We also synthesized these findings into a decision model (Figure 6) with six interaction loops, providing a grounded framework for understanding when users expect intervention support, when they prefer autonomous execution, and how agents can anticipate and adapt to these fluid transitions.
Drawing from these findings, we derive design implications for adaptive collaborative agents, examine concurrent interactions as behavioral signals for workstyle learning, and discuss how these principles extend beyond creative domains. Our findings position concurrent interaction not as a replacement for delegation, but as what makes delegation work better. By building mutual visibility and understanding that allows agents to learn how users work, full delegation becomes not a leap of faith, but a natural extension of an already-established collaboration. We also release our codes with a dataset of 214 interaction turns to support future research on human-agent co-creative collaboration by simulation. Our contributions include:
@@
  • A taxonomy of user action patterns in human-agent interaction, comprising five categories and ten codes characterizing how users engage with agents during execution.
@@
  • Design implications for adaptive collaborative agents, covering process visibility adaptation, behavioral signal interpretation, and proactive execution management.
  • A dataset of 214 human-agent co-creative interaction turns, uniquely capturing concurrent interaction logs (when and how users work alongside agents)
@@ 2.1. Dynamics of Flexible Collaboration: From Turn-taking to Synchrony
Human collaboration is a dynamic spectrum of interaction modes (Tang, 1991). While turn-taking reduces cognitive load and conflict in complex tasks through sequential actions (Kirschner et al., 2009; Sacks et al., 1974), creative domains often demand synchronous collaboration (Nam and Sakong, 2009). Creative work involves the co-evolution of problem and solution, where design ideas are refined through immediate visual feedback and spontaneous mutual inspiration (Maher and Poon, 1996; Ishii et al., 1993). Therefore, partners often bump in for immediate adjustments or perform simultaneous edits to interleave workflows, ensuring that the evolving design remains a shared mental model (Dourish and Bellotti, 1992; Tang, 1991).
@@
Modern tools like Figma and Google Slides support this fluidity by fostering workspace awareness. This is enabled through visual cues like telepointers (moving cursors) and object highlighting, allowing users to anticipate moves and intervene without explicit communication (Gutwin and Greenberg, 2002; Gutwin et al., 1996). However, most co-creative AI systems remain confined to a rigid wait-your-turn paradigm, largely limiting human-AI interaction to a prompt-and-output cycle, lacking the mutual awareness essential for true creative partnership. Moreover, working with generative AI has been shown to reduce divergent thinking (Anderson et al., 2024), underscoring that how humans and AI collaborate matters as much as whether they do.
2.2. Paradigm Shifts in Human-Agent Collaboration
@@
AI is evolving from a command-based tool into a proactive “co-worker” that actively participates within shared workspaces (Horvitz, 1999). This shift has been driven by frameworks enabling models to leverage diverse tools (Schick et al., 2023) and interleave reasoning with actions (Yao et al., 2023), enabling agents for programming (by Anthropic, ; Jimenez et al., 2023; Sapkota et al., 2025), web automation (AI, ; Deng et al., 2023; Jiang and Xia, 2025), visual design (Make, ), research (Gottweis et al., 2025; Feng et al., 2024), and data analysis (Hong et al., 2025). As these agents expose their reasoning and intermediate actions, this affords greater interpretability by allowing users to monitor agent behavior in real time (Amershi et al., 2019; Prasongpongchai et al., 2025).
@@
Yet most collaboration between human users and AI agents remains rooted in turn-taking. This limits users’ ability to correct agents mid-execution (Amershi et al., 2019) to prevent misalignments as the task progresses. In this work, we explore the potential of concurrent human-agent collaboration not as a replacement for turn-taking, but as an additional mode—enabling agents to move beyond delegated executors to responsive partners capable of accommodating spontaneous interventions within a shared workspace. This reflects a broader recognition that full delegation—optimizing purely for efficiency—risks devaluing the creative process itself (Almeda et al., 2025).
@@ 2.3. Design Agents: From Workflow Automation to Process-Visible Collaboration
Existing agents for supporting the design process can be categorized into three distinct approaches based on how they structure their tasks and interact with users. The first category is subtask design agents, which support specific stages within an existing design workflow by automating specific and tedious subtasks, adding details (Lawton et al., 2023; Karimi et al., 2020; Zhou et al., 2024) or generating alternatives (Make, ), while the overall workflow remains human-driven. The second, process-structured agents, operate through predefined node-based workflows (Hao et al., 2025; Wang et al., 2025; Ding et al., 2023; You et al., 2025), supporting guided co-creation, but constrained by predefined logical nodes. The third, output-centric agents, expose only a textual reasoning trace before delivering final results (Stitch, ; Lovart, ; Khurana et al., 2025).
Across all three, the agent’s execution remains hidden from the user’s workspace: users interact only through requests and final outputs, with no visibility into work in progress. While agents in other domains (e.g., programming) increasingly expose execution steps through direct tool operations, design tools have yet to explore how “live” tool-use—such as visible cursor movements and object manipulation—could reshape human-agent collaboration. This suggests an open question of whether process visibility could enable more diverse collaboration modes beyond simple turn-taking.
@@ 3. Study 1: Initial Technology Probe Study
To explore how process visibility creates opportunities for co-creative interactions and reveals the coordination challenges it entails in creative workflow, we implemented an initial design probe and conducted an exploratory study with 10 designers.
@@ 3.1.1. Three Basic Components for Collaborative Agent Design
Existing design agents are largely opaque—they accept instructions and return text reasoning (or high-level sub-goals) and final outputs, giving users little visibility into execution. To make agent execution observable and collaborative, we grounded the probe’s design consideration in three core components drawn from human-human collaboration literature, where these principles have been shown to support coordination, mutual understanding, and fluid collaboration, reasoning that they could equally scaffold more natural human-agent collaboration: (1) Interaction Modality for peer-like communication, (2) Process Transparency for observable agent work, and (3) Presence Awareness for mutual attention visualization.
(1) Interaction Modality: Voice Interaction with Flexible Object Referencing. The first probe uses voice as the primary interaction channel, enabling natural turn-taking, real-time intention sharing, and a sense of horizontal partnership (Hutchins, 1995; Reicherts et al., 2022; Bickmore and Picard, 2005). Users can also reference canvas objects directly during speech, enabling fluid deictic communication (e.g., “make this larger” while clicking).
(2) Process Transparency: Sharing Implementation Workflow. The first probe system exposes its design process in real-time through step-by-step execution, so users observe not just what the agent produces but how it progresses, establishing common ground around both goals and process (Robertson, 2002; Cila, 2022; Hill et al., 1992).
(3) Presence Awareness: Visualizing Attention in the Shared Workspace Following Gutwin’s workspace awareness framework (Gutwin and Greenberg, 2002), the agent’s position reveals its current focus, enabling users to anticipate upcoming actions and coordinate accordingly (Dourish and Bellotti, 1992; Fussell et al., 2000).
@@ 3.1.2. System Design
Figure 2. Interaction flow of the first probe system. (a) The user sends a voice message (b) while selecting canvas elements for context. (c) The agent generates an action plan based on the voice transcription with selections and iterates through ReAct cycles. (d-f) Each iteration’s progress is reflected immediately on the canvas as tools execute during the act stage.
@@
Interface and Interaction Scenario. The first probe is implemented as a Figma plugin where users combine voice input with canvas element selection to provide canvas context-based instructions to the agent (Figure 2-a-b). The agent responds through iterative tool execution—reasoning, acting, and self-evaluating in cycles—with each step reflected immediately on the canvas via a floating agent character alongside interim messages. (Figure 2-c-f).
Agent Design and Pipeline. The first probe agent is built on the ReAct structure (Yao et al., 2023), operating through a three-step iteration cycle: (1) reason, (2) act, and (3) feedback (Figure 2-c; Appendix A.4). At the reasoning stage, the agent selects from 38 Figma operations (e.g., create rectangle) and generates an interim progress message, which is executed on the canvas alongside the agent character. After each cycle, the summary module generates a text description of what occurred, while the feedback module compares the canvas state before and after execution to evaluate plan completion. If the plan is fulfilled, the agent terminates and delivers a final response; otherwise, it begins a new iteration, up to a maximum of 10 cycles .
We use Claude Sonnet 4.5 (Anthropic, 2026b) as the main reasoning model and Claude Haiku 4.5 (Anthropic, 2026a) for all other modules. Full prompt instructions, including Figma design guidelines (adapted from (Jeong et al., 2026)) and the complete set of 38 Figma operation tools (See Appendix A.1.2).